OpenStudy (anonymous):

solve the equation by completing the square. Round to the nearest hundredth if necessary. x2 + 7x = –12 (1 point) 3, –4 –3, –4 3, 4 –3, 4

OpenStudy (anonymous):

x^2 + 7x = –12

OpenStudy (anonymous):

well you know the two x are 3 and 4. since both 7 and 12 have + , you know that 3 and 4 are also positive

OpenStudy (anonymous):

x=4; -3

OpenStudy (anonymous):

\[x ^{2}+4x+3x+12\] x(x+4)+3(x+4) x+4=0; x+3=0 x=-4; -3